The American Heart Association is about to conduct an anti-smoking campaign and wants to know the fraction of Americans over 41 who smoke.

Suppose a sample of 830 Americans over 41 is drawn. Of these people, 673 don't smoke. Using the data, construct the 85% confidence interval for the population proportion of Americans over 41 who smoke. Round your answers to three decimal places.
Answer #1

T(who dont smoke) = 673

T(who smoke) = 830-673 = 157

Normal approximation to the binomial calculation:

Standard error of the mean = SEM = ?x(N-x)/N3 = 0.014

? = (1-CL)/2 = 0.075

Standard normal deviate for ? = Z? = 1.440

Proportion of positive results = P = x/N = 0.189

Lower bound = P - (Z?*SEM) = 0.170

Upper bound = P + (Z?*SEM) = 0.209